A current of 4.90 A is passed through a Fe(NO3)2 solution for 1.10 hours. How much iron is plated out of the solution?
Answer #1

Fe+2 + 2e- --> fe(s)

Total C = I*t = 4.9*1.1*3600 = 19404 C

since

1 mol of e- = 96500 C

then

19404/96500 = 0.20107 mol of e-

mol of Fe = 0.20107/2 = 0.100535 mol of Fe

mass = mol*MW = 0.100535*55.5 = 5.529425 g of Fe